# Identification of phase correlations in Financial Stock Market Turbulence

### v1

@@ E.2 Extended Fourier Transform
Here we multiply two different Fourier components $F(\omega_{\alpha})$ and $F(\omega_{\beta})$ and then multiply the result with the complex conjugate of Fourier modes at frequency $(\omega_{\alpha}+\omega_{\beta})$ as previously mentioned in Sharma et al. (2025).
@@ E.3 Test of EFT on pure noise
using TARA computational framework described in Mukherjee and Ganesh (2019), Saikia and Mukherjee (2024), Mukherjee et al. (2019b). Here we use a spatial grid resolution of $N=2^{10}$ and temporal resolution of $dt=10^{-4}$. The simulation box is of length $L=2\pi$ and the initial condition is set as $u(0)=\sin(x)$. We choose $\nu=3\times 10^{-3}$ and the amplitude of forcing is set to $A=6$. The external forcing is applied by a white noise with amplitude $A$ and zero mean. Since this represents an externally forced decaying turbulence, the energy spectra in large wavenumbers show a slight depreciation, as expected from any standard turbulence spectrum (see Figure 5).

### v2

@@ E.2 Extended Fourier Transform
Here we multiply two different Fourier components $F(\omega_{\alpha})$ and $F(\omega_{\beta})$ and then multiply the result with the complex conjugate of Fourier modes at frequency $(\omega_{\alpha}+\omega_{\beta})$ as previously mentioned in [54].
@@ E.3 Test of EFT on pure noise
using TARA computational framework described in [37], [55], [56]. Here we use a spatial grid resolution of $N=2^{10}$ and temporal resolution of $dt=10^{-4}$. The simulation box is of length $L=2\pi$ and the initial condition is set as $u(0)=\sin(x)$. We choose $\nu=3\times 10^{-3}$ and the amplitude of forcing is set to $A=6$. The external forcing is applied by a white noise with amplitude $A$ and zero mean. Since this represents an externally forced decaying turbulence, the energy spectra in large wavenumbers show a slight depreciation, as expected from any standard turbulence spectrum (see Figure 5).